OpenStudy (anonymous):

Can someone help me on how to solve this? cos[Arc tan (5/12) - Arc cot(4/3) ]

OpenStudy (anonymous):

me.. use pythagorean theorem...

OpenStudy (anonymous):

like how?

OpenStudy (anonymous):

for example ... cos ( arc sin 1/2 ) .. how to solve that??

OpenStudy (anonymous):

√3 /2

OpenStudy (anonymous):

correct.. then, you will solve that by using that...

OpenStudy (anonymous):

aw ok ok thank yew ^^

OpenStudy (anonymous):

but, I'm not sure.. haha, sorry. :(

OpenStudy (tkhunny):

Of course, you may wish to expand it, first. cos(x-y) = cos(x)(y)+sin(x)sin(y)

OpenStudy (anonymous):

that would be cos(√3/2 -4/5) ?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

then what? :D

OpenStudy (anonymous):

8/65? hm, am i right, @tkhunny ?

OpenStudy (tkhunny):

Absolutely not. \((a + b)^{2} \ne a^{2} + b^{2}\) generally. \(log(a + b) \ne log(a) + log(b)\) generally. \(\cos(a + b) \ne \cos(a) + \cos(b)\) generally. Start here.